OpenStudy (anonymous):

WILL MEDAL AND FAN IF HELP!!!!!! NEED HELP BAD!!!!D: PLEASE!!!!!!!! WILL YOU PLEASE ANSWER THESE FIVE FOR ME????????

OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations?

OpenStudy (anonymous):

2x+y=-8 x-y=-4

OpenStudy (anonymous):

A. (−3, −2) B. (−1, 3) C. (1, 5) D. (−4, 0)

OpenStudy (anonymous):

any one?

OpenStudy (ccswims):

D

OpenStudy (anonymous):

Yes hold on I am working out the problem

OpenStudy (ccswims):

2 x -4 + 0 = -8

OpenStudy (ccswims):

and -4 + 0 = -4

OpenStudy (anonymous):

D is correct.

OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? x-4y=6 3x+y=-8 A. (2, −1) B. (1, −9) C. (−2, −2) D. (−6, −3)

OpenStudy (ccswims):

C

OpenStudy (anonymous):

Which ordered pair is the solution to the system of equations? -5x-3y=12 x-3y=-6 A. (3, −3) B. (−3, 1) C. (6, 4) D. (9, 5)

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

your a proffesional @ccswims

OpenStudy (anonymous):

any idea?

OpenStudy (ccswims):

B

OpenStudy (anonymous):

2 more

OpenStudy (ccswims):

ok

OpenStudy (anonymous):

Tickets to the zoo cost $14 for adults and $10 for children. A group 12 people went to the zoo, and the tickets cost $140. The system of equations models this situation, where x is the number of adults and y is the number of children. x+y=12 14x+10y=140 How many adults and how many children were in the group? A. 5 adults and 7 children B. 6 adults and 6 children C. 7 adults and 5 children D. 8 adults and 4 children

OpenStudy (ccswims):

A

OpenStudy (anonymous):

LLAAAAAAASSSSSTTTT OOONNNNEEE

OpenStudy (ccswims):

ok

OpenStudy (anonymous):

A container holds a combination of nickels and quarters. There are 156 coins in all, and their total value is $21.60. The system of equations models this situation, where x is the number of nickels and y is the number of quarters. x+y=156 0.05x+0.25y=21.6 How many nickels and quarters are in the container? A. 118 nickels and 38 quarters B. 106 nickels and 50 quarters C. 87 nickels and 69 quarters D. 72 nickels and 84 quarters

OpenStudy (ccswims):

C

OpenStudy (anonymous):

i love i love you i love you

OpenStudy (anonymous):

YOU MADE ME PASS NOW I JUST NEED TO SEE WHAT I NEED TO DO NOW

OpenStudy (anonymous):

you are my best friend

OpenStudy (anonymous):

can you help me with 5 more?

OpenStudy (anonymous):

then ill leave you alone

OpenStudy (ccswims):

yea sure

OpenStudy (anonymous):

For the graph shown, select the statement that best represents the given system of equations

OpenStudy (anonymous):

hold on theres an attatchment

OpenStudy (anonymous):

2y-x=2 x-2y=8 A. inconsistent B. not enough information C. coincident D. consistent and independent

OpenStudy (ccswims):

I believe it is A

OpenStudy (anonymous):

i think im done for hte day thanks lots can you help me with tomorrow? what time will you be on?

OpenStudy (ccswims):

Basically anytime in the morning that isn't 9:30 - 9:50 and I won't be on at 12:30 - 4, and anytime after 4 is ok